# Visual Evolutionary Optimization on Graph-Structured Combinatorial Problems with MLLMs: A Case Study of Influence Maximization

### v2

@@ Abstract
Graph-structured combinatorial problems in complex networks are prevalent in many domains, and are computationally demanding due to their complexity and non-linear nature. Traditional evolutionary algorithms (EAs), while robust, often face obstacles due to content-shallow encoding limitations and lack of structural awareness, necessitating hand-crafted modifications for effective application. In this work, we introduce an original framework, visual evolutionary ptimization (VEO), leveraging multimodal large language models (MLLMs) as the backbone evolutionary optimizer in this context. Specifically, we propose a context-aware encoding way, representing the solution of the network as an image. In this manner, we can utilize MLLMs’ image processing capabilities to intuitively comprehend network configurations, thus enabling machines to solve these problems in a human-like way. We have developed MLLM-based operators tailored for various evolutionary optimization stages, including initialization, crossover, and mutation. Furthermore, we propose that graph sparsification can effectively enhance the applicability and scalability of VEO on large-scale networks, owing to the scale-free nature of real-world networks. We demonstrate the effectiveness of our method using a well-known task in complex networks, influence maximization, and validate it on eight different real-world networks of various structures. The results have confirmed VEO’s reliability and enhanced effectiveness compared to traditional evolutionary optimization.
@@ I Introduction
  • Generalization and Adaptability: MLLMs can adapt across diverse and dynamic graph structures without requiring problem-specific algorithm design. They can also flexibly respond to changes in topology, offering broad applicability and real-time adaptability.
@@ III-D MLLM-based Evolutionary Optimizer
Algorithm 3 Visual Evolutionary Optimization using MLLMs
@@ IV-C Examination of MLLM as Initialization Operator
TABLE IV: Statistical analysis of fitness values obtained by different population initialization modes across eight networks.
@@
To compare different methods, we use the Wilcoxon test with a 95% confidence interval to statistically examine their difference. The ‘+’ indicates statistically better performance, $\approx$’ indicates no significant difference, and ‘-’ indicates statistically worse performance. Table IV-C presents the optimization results regarding fitness values achieved through different initialization modes across eight networks. The fitness values are reported as averages with standard deviations. The MLLM-based method consistently outperforms other modes across all networks, as indicated by the highest fitness values and its average ranking of 1.00, demonstrating its superiority. This analysis highlights the superior capability of MLLMs in initializing the population for evolutionary optimization more effectively compared to traditional approaches. In addition, the results show that Refined Random initialization (sampled from the simplified network) outperforms the other three non-MLLM initialization methods (derived from the original network) in 7 out of 8 cases except for Netscience, demonstrating the effectiveness of our graph sparsification approach.
@@
TABLE VI: Statistical analysis of fitness values obtained by different reproduction modes across eight networks, using MLLM-driven evolutionary operations. The networks are visualized in KK and FR styles, and the MLLM operates under One-Shot and Two-Phase mutation strategies.
@@ IV-D Ablation Study of MLLM as Reproduction Operator
Furthermore, the results across networks plotted in both KK and FR styles indicate that the KK layout style generally achieves marginally better fitness results compared to those in the FR style. While most cases exhibit no statistically significant differences, the observed variations in certain instances indicate that the graph layout might be the critical factor affecting the performance of MLLM-driven optimization.
Figure 3 presents the optimization landscape of VEO against traditional evolutionary optimization across eight networks. The VEO approach, facilitated by MLLMs, consistently outperforms the traditional method, achieving higher fitness scores and demonstrating accelerated improvements in the initial iterations. The shaded regions indicate variability in the results, showing the robustness and reliability of the VEO process. These findings suggest that MLLMs provide insightful guidance, facilitating optimization in graph-structured combinatorial problems and achieving superior outcomes with fewer iterations.
Figure 3: Fitness evolution over iterations for different reproduction modes across eight different networks. Shaded areas indicate standard deviations over multiple runs.
@@ IV-F Foundation Model Comparison
Figure 6: Fitness evolution over iterations for MLLM-based reproduction strategies using different models across eight different networks. Shaded areas indicate standard deviations over multiple runs.

### v4

@@ Abstract
Graph-structured combinatorial problems in complex networks are prevalent in many domains, and are computationally demanding due to their complexity and non-linear nature. Traditional evolutionary algorithms (EAs), while robust, often face obstacles due to content-shallow encoding limitations and lack of structural awareness, necessitating hand-crafted modifications for effective application. In this work, we introduce an original framework, visual evolutionary optimization (VEO), leveraging multimodal large language models (MLLMs) as the backbone evolutionary optimizer in this context. Specifically, we propose a context-aware encoding scheme, representing the solution of the network as an image. In this manner, we can utilize MLLMs’ image processing capabilities to intuitively comprehend network configurations, thus enabling machines to solve these problems in a human-like way. We develop MLLM-based operators tailored for various evolutionary optimization stages, including initialization, crossover, and mutation. Furthermore, we propose that graph sparsification can effectively enhance the applicability and scalability of VEO on large-scale networks, owing to the scale-free nature of real-world networks. We demonstrate the effectiveness of our method using a well-known task in complex networks, influence maximization, and validate it on eight different real-world networks of various structures. The results confirm VEO’s reliability and enhanced effectiveness compared to traditional evolutionary optimization.
@@ I Introduction
  • Generalizability and Adaptability: MLLMs can adapt across diverse and dynamic graph structures without requiring problem-specific algorithm design. They can also flexibly respond to changes in topology, offering broad applicability and real-time adaptability.
@@ III-D MLLM-based Evolutionary Optimizer
Algorithm 3 Visual Evolutionary Optimization
@@ IV-C Examination of MLLM as Initialization Operator
TABLE IV: Statistical analysis (Mean and Standard Deviation (SD)) of fitness values obtained by different population initialization modes across eight networks. The optimization follows the normal evolutionary optimization and the population evolves for 50 generations.
@@
To compare different methods, we use the Wilcoxon test with a 95% confidence interval to statistically examine their difference. The ‘+’ indicates statistically better performance, $\approx$’ indicates no significant difference, and ‘-’ indicates statistically worse performance. Table IV-C presents the optimization results regarding fitness values achieved through different initialization modes across eight networks. The fitness values are reported as averages with standard deviations (SD). The MLLM-based method consistently outperforms other modes across all networks, as indicated by the highest fitness values and its average ranking of 1.00, demonstrating its superiority. This analysis highlights the superior capability of MLLMs in initializing the population for evolutionary optimization more effectively compared to traditional approaches. In addition, the results show that Refined Random initialization (sampled from the simplified network) outperforms the other three non-MLLM initialization methods (derived from the original network) in 7 out of 8 cases except for Netscience, demonstrating the effectiveness of our graph sparsification approach.
@@
TABLE VI: Statistical analysis (Mean and Standard Deviation (SD)) of fitness values obtained by different reproduction modes across eight networks, using MLLM-driven evolutionary operations. The networks are visualized in $K$ and FR styles, and the MLLM operates under One-Shot and Two-Phase mutation strategies. The population is initialized using the refined random strategy within the sparsified network domain and then evolves for 10 generations.
@@ IV-D Ablation Study of MLLM as Reproduction Operator
Furthermore, the results across networks plotted in both KK (Kamada–Kawai) and FR (Fruchterman–Reingold) styles indicate that the KK layout style generally achieves marginally better fitness results compared to those in the FR style. While most cases exhibit no statistically significant differences, the observed variations in certain instances indicate that the graph layout might be the critical factor affecting the performance of MLLM-driven optimization.
Figure 3 presents the optimization landscape of VEO against traditional evolutionary optimization across eight networks. The VEO approach, facilitated by MLLMs, consistently outperforms the traditional method, achieving higher fitness scores and demonstrating accelerated improvements in the initial iterations. The shaded regions indicate variability in the results, showing the robustness and reliability of the VEO. These findings suggest that MLLMs provide insightful guidance, facilitating optimization in graph-structured combinatorial problems and achieving superior outcomes with fewer iterations.
Figure 3: Fitness evolution over iterations for different reproduction modes across eight different networks. Shaded areas indicate Standard Error of the Mean (SEM) over multiple runs.
@@ IV-F Foundation Model Comparison
Figure 6: Fitness evolution over iterations for MLLM-based reproduction strategies using different models across eight different networks. Shaded areas indicate Standard Error of the Mean (SEM) over multiple runs.